a sample contains of 14.2g chlorine gas.What is the volume of chlorine gas at room temperature and pressure? [Relative atomic mass,Cl=35.5;Molar volume of gas=24dmmol-1 at room temperature and pressure


1
Expert's answer
2021-10-01T12:19:02-0400

M(Cl2) = 71 g/mol

n(Cl2) "= \frac{m}{M} = \frac{14.2}{71} = 0.2 \;mol"

V(Cl2) "= 24 \times 0.2 = 4.8 \;dm^3"

Answer: 4.8 dm3